A man invests $1200, part at 5% and part at 4% per annum, and his total income is $55.50. How much did he invest at each rate?
Answer by mananth(16949)

Part I 5.00% per annum ------------- Amount invested =x
Part II 4.00% per annum ------------ Amount invested = y
1200
Interest----- 55.50

Part I 5.00% per annum ---x *
Part II 4.00% per annum ---y
Total investment
x + 1 y= 1200 -------------1
Interest on both investments
5.00% x + 4.00% y= 55.5
Multiply by 100
5 x + 4 y= 5550.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -6000.00
Add this to (2)
0 x -1 y= -450
divide by -1
y = 450 - x)
Part I 5.00% $ 750
Part II 4.00% $ 450

CHECK
750 --------- 5.00% ------- 37.50
450 ------------- 4.00% ------- 18.00
Total -------------------- 55.50